OpenStudy (anonymous):

Michael and Minnie are gathering coats for a coat-drive at their school. Minnie gathers 3 times as many coats as Michael and together they have 52 coats. How many coats did Michael contribute?

OpenStudy (ddcamp):

Let A be the number of coats Michael brought, and B be the number that Minnie brought. The total that they brought together would be A+B = 52 Since Minnie brought three times as many coats as Michael, we know that B=3A Can you solve the system for A?

OpenStudy (anonymous):

I don't know it :/

OpenStudy (anonymous):

wait i got it

OpenStudy (ddcamp):

We have two equations, and we know that A and B are the same in both equations: A + B = 52 B = 3A Since the B in the first equation is the same as the B in the second equation, we can substitute one value of B for the other. Since B=3A, we can change the first equation to say: A + B = 52 A + (3A) = 52 Can you solve that for A?